# Can Human Drivers and Connected Autonomous Vehicles Co-exist in Lane-Free Traffic? A Microscopic Simulation Perspective

### v1

@@ Abstract
Recent advancements in connected autonomous vehicle (CAV) technology have sparked growing research interest in lane-free traffic (LFT). LFT envisions a scenario where all vehicles are CAVs, coordinating their movements without lanes to achieve smoother traffic flow and higher road capacity. This potentially reduces congestion without building new infrastructure. However, the transition phase will likely involve non-connected actors such as human-driven vehicles (HDVs) or independent AVs sharing the roads. This raises the question of how LFT performance is impacted when not all vehicles are CAVs, as these non-connected vehicles may prioritize their own benefits over system-wide improvements. This paper addresses this question through microscopic simulation on a ring road, where CAVs follow the potential lines (PL) controller for LFT, while HDVs adhere to a strip-based car-following model. The PL controller is also modified for safe velocities to prevent collisions. The results reveal that even a small percentage of HDVs can significantly disrupt LFT flow: 5% HDVs can reduce LFT’s maximum road capacity by 16%, and a 20% HDVs nearly halves it. The study also develops an adaptive potential (APL) controller that forms APL corridors with modified PLs in the surroundings of HDVs. APL shows a peak traffic flow improvement of 23.6% over the PL controller. The study indicates that a penetration rate of approximately 60% CAVs in LFT is required before significant benefits of LFT start appearing compared to a scenario with all HDVs. These findings open a new research direction on minimizing the adverse effects of non-connected vehicles on LFT.
@@
  • The LFT traffic flow via Potential Lines controller halves by 20% HDV penetration rate
@@
  • A peak improvement of 23.6% was observed via the APL strategy over PL controller
@@ 1 Introduction
The LFT was inspired by the vehicles driving without strict lane discipline in some countries, referred to as lane-less traffic Sekeran et al. (2022); Papageorgiou et al. (2021). This allows them to utilize the whole width of the road; however, with a significantly higher risk of accidents Mukherjee and Mitra (2022). Apart from automation that allows fast and accurate decisions in LFT, the main difference between the two modes is that while in lane-less traffic, communication with other drivers is limited to honking, headlight flashing, or hand gesturing, the vehicles in LFT can fully communicate their intended trajectories and other important information to significantly larger areas. This allows LFT vehicles to coordinate their movements in a way that substantially improves the system-wide maximum flow rate compared to lane-based traffic for the same road width Rostami-Shahrbabaki et al. (2023b). However, the exact flow rate improvement depends on the LFT controller used.
While the LFT strategy has a high potential to improve traffic conditions, its realization may still be decades ahead. First, significant technological developments are required to ensure flawless inter-vehicular communication and reliable autonomous driving functionality. Second, the CAVs in LFT will be driving in a completely new traffic mode where safety will be of primary concern. Consequently, LFT must undergo rigorous testing before any CAV can be driven in LFT mode. However, unlike the functionality of current AVs, which can be tested with a single vehicle, the LFT would require tests with multiple CAVs operating in an unexplored driving mode. To solve this problem, some researchers have suggested using a driving simulator to study the safety as perceived by the traveler in LFT Sekeran et al. (2023). Nonetheless, it is expected that even if the cities are convinced of investing in the LFT, the transition phase or even the actual operation of LFT may still involve some traffic participants who do not necessarily coordinate their movements with other vehicles. This can happen for multiple reasons, for example, the involvement of human drivers, having an older AV without a communication module, or a CAV temporarily losing its communication capability. Under all of these circumstances, the question remains: how will the efficiency of the LFT be affected by it? It is equally important to ask how much the proportion of CAVs should be to achieve the benefits of LFT.
@@
To study the impact of the whole setup, microscopic simulations of a 1 $km$ ring road are used. The human drivers are modeled using the car-following model by Mathew et al. (2015), designed to simulate lane-less traffic. For the CAVs, the paper uses the potential line (PL) controller introduced by Rostami-Shahrbabaki et al. (2023b). Inspired by the human model of Mathew et al. (2015), the paper also introduces the concept of safe acceleration into the PL controller as a secondary measure to avoid any potential collision. Finally, an adaptive potential lines (APL) controller is introduced that improves the flow of CAVs mixed with HDVs in LFT. The new controller forms APL corridors with modified PLs in the vicinity of HDVs.
@@
1.1 Background
@@
Since LFT allows for lateral freedom of CAVs, compared to lane-based traffic, novel driving strategies have been proposed for vehicle navigation in this new environment. A common approach comprises defining artificial potential fields around each vehicle for collision avoidance Malekzadeh et al. (2022), as shown in Figure 2(a), and including additional controllers for achieving other objectives such as driving close to the desired speed, staying within the road boundary, and driving energy-efficiently. Yanumula et al. (2023) formulated all the objectives in an optimal control problem and solved it for each vehicle in real-time. A nonlinear model predictive control (NMPC) approach is proposed in Levy and Haddad (2022) for cooperative driving among autonomous vehicles. A more structured lane-free traffic is proposed in Zhang et al. (2023). In this approach, an artificial potential line (PL) is assigned for each vehicle as the desired lateral location, as illustrated in 2(b). This led to more laminar traffic flow, eliminating unnecessary lateral movement of vehicles. Other approaches, such as nonlinear controllers Theodosis et al. (2023), have also been proposed that use a more complex dynamic model of vehicles. While the majority of LFT research focuses on freeway networks, few have considered elements of urban networks such as beltways Rostami-Shahrbabaki et al. (2023b), intersections Stüger et al. (2023); Malcolm and Bogenberger (2025), and roundabouts Naderi et al. (2024). In addition to conventional control approaches, reinforcement learning-based approaches have also been used, showcasing comparable benefits for driving comfort and traffic efficiency Berahman et al. (2024, 2022); Karalakou et al. (2022, 2023). In addition, LFT calls for novel ideas such as vehicle flocking Rostami-Shahrbabaki et al. (2023a, 2022) or snake-like platooning Dabestani et al. (2024).
@@ 2.2.1 Longitudinal Movement
In the current study, after determining the $v^{safe}_{x,i}$, the difference to the current speed $v^{diff}_{i}=v_{safe}-v_{x,i}$ is calculated, which is then used to calculate the safe acceleration:
[…]
where $A^{+}_{x}$ is the vehicle’s maximum acceleration ability. Eq. 2 shows that a vehicle’s acceleration and deceleration are constrained by its abilities, i.e. $A^{-}_{x}$ and $A^{+}_{x}$, respectively.
@@ 2.2.2 Lateral Movement
The lateral movements in the human model of Mathew et al. (2015) are also based on strips rather than lanes. The decision to change a strip is dependent on the benefit of changing the strip, measured in terms of speed gain. A vehicle is allowed to change only one strip in a time step; however, the model considers multiple strips in calculating the benefits since the driver’s overtaking maneuver may require changing multiple strips. Thus, the decision to change the strip involves multiple steps. First, for a vehicle $i$, the benefit of changing the strip from a currently occupied strip $s_{c}$ to a destination strip $s_{d}$ is calculated using the following formula:
@@ 2.3 Potential Lines Controller for LFT
The following details the components used for the PL controller:
@@ 2.3.1 PL based Lateral Control
$f_{pl,i}=K_{plf}\cdot(y_{pl,i}-y_{i})$ (5)
where $K_{plf}$ is the controller gain for PL force. In the current study, it is set to 0.12.
@@ 2.3.2 Cruise Controller based Longitudinal Control
[…]
where $A^{+}_{x}$ is the acceleration ability. Rostami-Shahrbabaki et al. (2023b) did not use $A^{+}_{x}$ in their formulation. However, this study adds it to also cater for the situation when the vehicle comes to a complete halt. The current study uses $K_{px}$ value of 1.0, which makes the $f^{cc}_{x,i}$ to be directly determined by $A^{+}_{x}$ and $v_{des,i}$, in consistency with the human model (Eq. 2).
@@ 2.3.5 Extension of Collision Avoidance via Safe Acceleration
In view of the above, the current study introduces the constraint of safe acceleration into LFT controller. The first step in this process is selecting a leader vehicle. Similar to the human model, the vehicle with the shortest longitudinal distance and an overlap in the lateral axis is chosen as the leader for vehicle $i$. However, unlike the human model, the overlap is directly calculated using vehicle widths and lateral positions without discretizing the lateral axis into strips. After determining the leader, the safe velocity $v^{safe}_{x,i}$ and the safe acceleration $a_{x,i}^{safe}$ are calculated using Eq. 1 and Eq. 2, respectively. $a_{x,i}^{safe}$ is then used to bound the longitudinal acceleration calculated by the PL controller; thus, the overall longitudinal acceleration is given as:

### v2

@@ Abstract
Recent advancements in connected autonomous vehicle (CAV) technology have sparked growing research interest in lane-free traffic (LFT). LFT envisions a scenario where all vehicles are CAVs, coordinating their movements without lanes to achieve smoother traffic flow and higher road capacity. This potentially reduces congestion without building new infrastructure. However, the transition phase will likely involve non-connected actors such as human-driven vehicles (HDVs) or independent AVs sharing the roads. This raises the question of how LFT performance is impacted when not all vehicles are CAVs, as these non-connected vehicles may prioritize their own benefits over system-wide improvements. This paper addresses this question through microscopic simulation on a ring road, where CAVs follow the potential lines (PL) controller for LFT, while HDVs adhere to a strip-based car-following model. The PL controller is also modified for safe velocities to prevent collisions. The results reveal that even a small percentage of HDVs can significantly disrupt LFT flow: 5% HDVs can reduce LFT’s maximum road capacity by 20% and a 40% HDVs nearly halves it, up until 100% HDVs where it drops by nearly 60%. The study also develops an adaptive potential line (APL) controller that forms APL corridors in the surroundings of HDVs. APL shows a peak traffic flow improvement of nearly 10% over the PL controller. The study indicates that a penetration rate of approximately 60% CAVs is required to start observing the major LFT benefits. These findings open a new research direction on minimizing the adverse effects of non-connected vehicles on LFT.
Comments on the current draft
This version corresponds to the final published article in Transportation Research Part C: Emerging Technologies (DOI: https://doi.org/10.1016/j.trc.2025.105315). It incorporates revisions made during peer review, including an improved literature review, clearer methodological descriptions, and explicit consideration of safety and comfort. These changes affected some simulation results (e.g., maximum road capacity in the all-CAV scenario, impacts of HDV penetration, and improvements via the Adaptive Potential Lines method). However, the main conclusions remain unchanged—for instance, a CAV penetration of 60% is still required before major advantages of LFT are observed. The above-linked article is published under the Creative Commons Attribution 4.0 International (CC BY 4.0) license (https://creativecommons.org/licenses/by/4.0/).
@@
  • The LFT traffic flow drops by 20% with just 5% HDV penetration rate
@@
  • A peak average improvement of nearly 10% was observed via the APL strategy over PL controller
@@ 1 Introduction
The LFT was inspired by the vehicles driving without strict lane discipline in some countries, referred to as lane-less traffic Sekeran et al. (2022); Papageorgiou et al. (2021). A key motivating factor was the significantly greater utilization of road width observed in lane-less traffic, as opposed to the roughly 50% occupancy typical in lane-based systems. This 50% estimate is based on a comparison between the average widths of vehicles (cars and trucks) and standard lane widthsPapageorgiou et al. (2021). While lane-less traffic enables higher road occupancy, it also introduces a substantially increased risk of accidents Mukherjee and Mitra (2022) . However, with the automation of vehicles in LFT, such risks can be mitigated through the integration of advanced safety features within the control algorithms. Apart from fast and accurate decisions in LFT, another major difference between lane-less and LFT traffic is that while in the former, communication with other drivers is limited to honking, headlight flashing, or hand gesturing, the vehicles in the latter can fully communicate their intended trajectories and other important information to significantly larger areas. This allows LFT vehicles to coordinate their movements in a way that substantially improves the system-wide maximum flow rate compared to lane-based traffic for the same road width Rostami-Shahrbabaki et al. (2023b). However, the exact flow rate improvement depends on the LFT controller used.
While the LFT strategy has a high potential to improve traffic conditions, its realization may still be decades away. First, significant technological developments are required to ensure flawless inter-vehicular communication and reliable autonomous driving functionality. Second, the CAVs in LFT will be driving in a completely new traffic mode where safety will be of primary concern. Consequently, LFT must undergo rigorous testing before any CAV can be driven in LFT mode. However, unlike the functionality of current AVs, which can be tested with a single vehicle, the LFT would require tests with multiple CAVs operating in an unexplored driving mode. To solve this problem, some researchers have suggested using a driving simulator to study the safety as perceived by the traveler in LFT Sekeran et al. (2023). Nonetheless, it is expected that even if the cities are convinced of investing in the LFT, the transition phase or even the actual operation of LFT may still involve some traffic participants who do not necessarily coordinate their movements with other vehicles. This can happen for multiple reasons, for example, the involvement of human drivers, having an older AV without a communication module, or a CAV temporarily losing its communication capability. Under all of these circumstances, the question remains: how will the efficiency of the LFT be affected by it? It is equally important to ask how much the proportion of CAVs should be to achieve the benefits of LFT.
@@
To study the impact of the whole setup, microscopic simulations of a 1 $km$ ring road are used. The simulation builds upon prior research, extending and refining earlier contributions. Figure 2 illustrates these foundational works and their relationship to the key contributions of the present study. The human drivers are modeled using the car-following model by Mathew et al. (2015), designed to simulate lane-less traffic. For the CAVs, the paper uses the potential line (PL) controller introduced by Rostami-Shahrbabaki et al. (2023b). Inspired by the human model of Mathew et al. (2015), the paper also introduces the concept of safe acceleration into the PL controller as a secondary measure to avoid any potential collision. Finally, an adaptive potential lines (APL) controller is introduced that improves the flow of CAVs mixed with HDVs in LFT. The new controller forms APL corridors with modified PLs in the vicinity of HDVs.
@@
Figure 2: The main contributions of the current work and its relation to the available literature.
2 Background
@@
Since LFT allows for lateral freedom of CAVs, compared to lane-based traffic, novel driving strategies have been proposed for vehicle navigation in this new environment Papageorgiou et al. (2024). A common approach comprises defining artificial potential fields around each vehicle for collision avoidance Malekzadeh et al. (2022), as shown in Figure 3(a), and including additional controllers for achieving other objectives such as driving close to the desired speed, staying within the road boundary, and driving energy-efficiently. Yanumula et al. (2023) formulated all the objectives in an optimal control problem and solved it for each vehicle in real-time. A more structured lane-free traffic is proposed in Zhang et al. (2023). In this approach, an artificial potential line (PL) is assigned for each vehicle as the desired lateral location, as illustrated in 3(b). This led to more laminar traffic flow, eliminating unnecessary lateral movement of vehicles. Other approaches, such as nonlinear controllers Theodosis et al. (2023), have also been proposed that use a more complex dynamic model of vehicles. While the majority of LFT research focuses on freeway networks, few have considered elements of urban networks such as beltways Rostami-Shahrbabaki et al. (2023b), intersections Stüger et al. (2023); Malcolm and Bogenberger (2025), and roundabouts Naderi et al. (2024). In addition to conventional control approaches, reinforcement learning-based approaches have also been used, showcasing comparable benefits for driving comfort and traffic efficiency Berahman et al. (2025, 2022); Karalakou et al. (2022, 2023). In addition, LFT calls for novel ideas such as vehicle flocking Rostami-Shahrbabaki et al. (2023a, 2022) or snake-like platooning Dabestani et al. (2024). Table LABEL:table:_literature_review presents an overview of the LFT control methods.
@@
Table 1: Overview of the literature related to LFT control methods.
Study
Environment
Setting
Traffic
Participants
LFT Control Method
Papageorgiou et al. (2021) Ring Road CAVs A conceptual work with target speed and potential field-based LFT control. According to the detailed version of the work Papageorgiou et al. (2019), the target speed controller uses Guass error function for cruise control and the potential field is based on $\Pi$ function.
Yanumula et al. (2021) Ring Road CAVs The work introduces an ellipsoidal potential field function combined. The LFT is formulated as a non-linear optimal control problem with multiple objectives and a model predictive control approach.
Malekzadeh et al. (2022) Ring Road CAVs Extends the controller of Papageorgiou et al. (2021). The forces are based on adaptive cruise control (ACC)-inspired logic using an Action Graph for collision avoidance and a boundary feedback controller to keep vehicles within the road edges.
Berahman et al. (2022) Ring Road CAVs Combines the ellipsoidal artificial potential field with deep reinforcement learning (RL) based LFT control.
Karalakou et al. (2022) Ring Road CAVs Test different reward functions for training an RL-based LFT control.
Rostami-Shahrbabaki et al. (2022) Ring Road CAVs A two-level controller for vehicle flocking is developed, where in the tactical layer, the control mode is defined, and in the operational layer, the individual and group movement of vehicles is controlled.
Yanumula et al. (2023) Ring Road CAVs It builds on top of Yanumula et al. (2021) by introducing vehicle nudging, improved obstacle avoidance, and adaptive desired speeds, enabling more efficient maneuvers in dense traffic. It also enhances the optimal control formulation with additional objectives for comfort and emergency collision handling.
Rostami-Shahrbabaki et al. (2023b) Ring Road CAVs Introduces the concept of assigning specific lateral positions, called potential lines, to vehicles according to their desired speeds. A Proportional controller is then used for lateral control of vehicles.
Zhang et al. (2023)
Motorway
with ramps
CAVs,
Emergency
Vehicle
The LFT control is the same as in Rostami-Shahrbabaki et al. (2023b). However, it uses the Probability Integral Transform for uniform lateral distribution of potential lines. An adaptation of the potential line strategy is used for emergency vehicle preemption.
Theodosis et al. (2023)
Ring Road,
Straight Road
CAVs It introduces a Lyapunov Function-based Cruise Controller based on a nonlinear bicycle dynamic model for vehicles.
Stüger et al. (2023) Intersection CAVs Intersection control using artificial potential fields and PID controller. The LFT control is divided over multiple modules such as: (1) module for flocking vehicles with similar routes (2) module for dynamic entry and exist points (3) module for reserving tiles for traversing the main intersection area.
Karalakou et al. (2023) Ring Road CAVs Extends the RL-based LFT controller of Karalakou et al. (2022) using multiple RL algorithms and improves the reward functions used for RL.
Rostami-Shahrbabaki et al. (2023a) Ring Road CAVs It proposed an algorithm for vehicular flocking based on (1) a Mexican hat function for flock formation, (2) a consensus algorithm for velocity matching, and (3) a navigation feedback.
Naderi et al. (2024)
Large
Roundabouts
CAVs Uses OD-specific movement corridors and space-dependent desired orientations to guide each vehicle. Non-linear feedback controllers are used for circular and straight movement, and linear feedback controllers are used for boundary control.
Berahman et al. (2025)
Ring Road,
Motorway with
ramps
CAVs Improves the control approach of Berahman et al. (2022) using a multi-agent and multi-task RL approach. The RL agents learn from shared experiences to achieve more complex driving tasks.
Dabestani et al. (2024) Straight Road CAVs It introduces a joint trajectory-based optimization approach forming 1-D snake-like or 2D deformable flocks. It minimizes several objectives like collision avoidance, deviation from desired speed, formation maintenance, etc.
Malcolm and Bogenberger (2025) Intersection CAVs, VRUs Uses a novel first come, first served reservation and trajectory planning approach called FERSTT (Fast Exploring Rule-based Spatiotemporal Trajectory Tree) for intersection control with vulnerable road users (VRUs). VRUs are given separate signal phases for crossing the intersection.
@@ 3.2.1 Longitudinal Movement
After determining the $v^{safe}_{x,i}$, the difference to the current speed $v^{diff}_{i}=v_{safe}-v_{x,i}$ is calculated, which is then used to calculate the safe acceleration:
$a_{x,i}^{safe}(k+1)=\begin{cases}min\big(\frac{v^{diff}_{i}}{\Delta T}\hskip 1.99997pt,\hskip 1.99997ptA^{-}_{x}\big)&\text{if }v^{diff}_{i}<0\\ min\big(\frac{v^{diff}_{i}}{\Delta T}\hskip 1.99997pt,\hskip 1.99997ptA^{+}_{x}\big)&\text{otherwise}\end{cases}$ (2)
where $A^{+}_{x}$ is the vehicle’s maximum acceleration ability. Eq. 2 shows that a vehicle’s acceleration and deceleration are constrained by its abilities, i.e., $A^{-}_{x}$ and $A^{+}_{x}$, respectively.
During experiments for the current study, it was noticed that setting $A^{-}_{x}$ and $A^{+}_{x}$ according to maximum vehicle capabilities leads to highly unrealistic, sudden vehicle movements with very high jerk. Thus, their values are set to a reasonable value range for comfortable acceleration and jerk Bae et al. (2019). Additionally, it was also noticed that after changing lateral positions, the new leader was sometimes not at a sufficient distance for vehicle $i$ to decelerate to the safe velocity without risking a collision, even when applying the desired deceleration $A^{-}_{x}$. To avoid this situation, if the braking distance is greater than the gap to the leader, i.e., $\frac{(v^{diff}_{i})^{2}}{2A^{-}_{x}}>g_{leader}(k)-g_{o}$, the HDV uses a higher deceleration value $A^{-}_{critical}$ to avoid collision.
@@ 3.2.2 Lateral Movement
The lateral movements in the human model of Mathew et al. (2015) are also based on strips rather than lanes. It is inspired by the lane change model of Ehmanns (2001), already implemented in SUMO Krajzewicz (2010). Mathew et al. (2015) applied this lane change approach to strips. The decision to change a strip is dependent on the benefit of changing the strip, measured in terms of speed gain. A vehicle is allowed to change only one strip in a time step; however, the model considers multiple strips in calculating the benefits since the driver’s overtaking maneuver may require changing multiple strips. Thus, the decision to change the strip involves multiple steps. First, for a vehicle $i$, the benefit of changing the strip from a currently occupied strip $s_{c}$ to a destination strip $s_{d}$ is calculated using the following formula:
@@ 3.3 Potential Lines Controller for LFT
At first glance, the PL controller might appear similar to the strip-based HDV model described in the previous section. However, key differences distinguish the two. In the HDV model, strips are used solely to discretize the lateral axis, enabling vehicles to occupy positions more refined than traditional lanes, but overall, it still follows a conventional car-following modeling approach. Additionally, the resolution of the strips constrains lateral acceleration to discrete values, as vehicles are restricted to a single strip at any given time and can shift only one strip per time step. In contrast, the PL controller does not rely on a car-following paradigm. Instead, it uses artificial potential fields to govern both lateral and longitudinal accelerations of CAVs. While these vehicles are encouraged to stay aligned with their assigned PLs, they are not strictly bound to them and can move freely along the lateral axis in response to the cumulative artificial forces. As a result, their lateral movements are significantly smoother than those of the HDV model.
With the above main distinguishing features, the following details the components used for the PL controller:
@@ 3.3.1 PL based Lateral Control
$f_{pl,i}=K_{pl}\cdot(y_{pl,i}-y_{i})-K_{pl,v}\cdot v_{y,i}$ (5)
where $K_{pl}$ and $K_{pl,v}$ are the controller gains for PL force.
@@ 3.3.2 Cruise Controller based Longitudinal Control
$v_{x,i}^{ts}=min\{v_{x,i}(k)+A^{+}_{x}\cdot\Delta T\hskip 1.99997pt,\hskip 1.99997ptv_{des,i}\}$ (7)
where $A^{+}_{x}$ is the preferred acceleration. Rostami-Shahrbabaki et al. (2023b) did not use $A^{+}_{x}$ in their formulation. However, this study adds it to also cater for the situation when the vehicle comes to a complete halt. The current study uses $K_{px}$ value of 1.0, which makes the $f^{cc}_{x,i}$ to be directly determined by $A^{+}_{x}$ and $v_{des,i}$, in consistency with the human model (Eq. 2).
@@ 3.3.4 Overall Accelerations
3.3.5 Boundary and other constraints for Lateral Acceleration
The lateral acceleration obtained in Eq. 10 must be restricted to keep the vehicles within the road boundary. A boundary feedback controller is used for this purpose to smoothly limit the lateral acceleration such that the vehicles do not cross the road boundaries Malekzadeh et al. (2022); Zhang et al. (2023). It is calculated as:
$a_{y,i}^{lim}=K_{b1}\cdot(\hat{y}-y_{i})-K_{b2}\cdot v_{y,i}$ (11)
where $\hat{y}$ is the road boundary limits that $y_{i}$ must not cross. Let $W_{i}$ represent the width of the vehicle $i$. Since $y_{i}$ is the lateral position of the center of the vehicle, setting the value of $\hat{y}$ to $Y_{l}-0.5W_{i}$ and $Y_{r}+0.5W_{i}$ determines the upper and lower limits for the lateral acceleration such that the vehicle does not cross the road boundary.
In addition to boundary control, the lateral acceleration is also bounded by minimum and maximum lateral acceleration ($A_{y}^{min}$ and $A_{y}^{max}$) and jerk ($J_{y}^{min}$ and $J_{y}^{max}$) limits for comfortable accelerations.
@@ 3.3.6 Extension of Collision Avoidance via Safe Acceleration
In view of the above, the current study introduces the constraint of safe acceleration into LFT controller. The first step in this process is selecting a leader vehicle. Similar to the human model, the vehicle with the shortest longitudinal distance and an overlap in the lateral axis is chosen as the leader for vehicle $i$. However, unlike the human model, the overlap is directly calculated using vehicle widths and lateral positions without discretizing the lateral axis into strips. After determining the leader, the safe velocity $v^{safe}_{x,i}$ and the safe acceleration $a_{x,i}^{safe}$ are calculated using Eq. 1 and Eq. 2, respectively. $a_{x,i}^{safe}$ is then used to bound the longitudinal acceleration calculated by the PL controller; thus, the longitudinal acceleration is given as:
@@
Finally, the $a_{x,i}(k+1)$ is also bounded by lower and upper acceleration ($A_{x}^{min}$ and $A_{x}^{max}$) and jerk ($J_{x}^{min}$ and $J_{x}^{max}$) limits, similar to lateral acceleration. It is important to note that even though the cruise controller in Eq. 7 uses the acceleration $A^{+}_{x}$ to achieve the desired speed, the actual acceleration of CAVs can be higher than $A^{+}_{x}$ due to combined nudging and repulsive forces of surrounding vehicles. This is in contrast to the HDV model, where no such forces exist to cause higher longitudinal acceleration than $A^{+}_{x}$. Thus, it is essential to bound the longitudinal acceleration of CAVs, which allows fine-tuning the rest of the LFT parameters for comfortable LFT movements.